# Directly Forecasting Belief for Reinforcement Learning with Delays

### v1

@@ 4.2 Multi-step Bootstrapping with DFBT
Then, we present our practical delayed RL method, named DFBT-SAC, by incorporating the trained DFBT $b_{\theta}$ with Soft Actor-Critic (Haarnoja et al., 2018). To improve the learning efficiency with the DFBT, the critic of DFBT-SAC is bootstrapped on the states predicted by the DFBT with the multi-step learning (Sutton & Barto, 2018; Hessel et al., 2018) and delay-free training techniques (Wu et al., 2024b; Kim et al., 2023). Specifically, given multi-step data $(x^{\text{tokens}}_{t},s_{t-\Delta+1:t})$, the critic $Q_{\psi}$ parameterized by $\psi$ is updated via:
@@ 5.2 Directly Forecasting Belief: Errors Analysis
Next, we theoretically show that our directly forecasting belief can effectively address the compounding errors. The sequence-modelling method estimates the states $s_{t-\Delta+1:t}$ based on the augmented state $x_{t}$ in parallel, instead of invoking $\Delta$ times iteratively, which effectively alleviates the source of compounding errors. Especially, the attention mechanism in the transformer can selectively capture the long-range relationships within the augmented state with long delays. We assume that the belief error between the directly forecasting belief and the ground-truth belief is bounded (5.6).
@@
The distance between the directly forecasting belief belief $b_{\theta}$ parameterized by $\theta$ and the ground-truth belief $b$ is bounded, it satisfies that $\forall x_{t}\in\mathcal{X}$, we have $\mathcal{W}(b(\cdot|{x_{t}})||b_{\theta}(\cdot|{x_{t}}))\leq\epsilon_{direct}$, where $\epsilon_{direct}:=\max_{i=1,\cdots,\Delta}\mathcal{W}(b^{(i)}(\cdot|{x_{t}})||b^{(i)}_{\theta}(\cdot|{x_{t}}))$.
@@
For the delay-free policy $\pi$ and the delayed policy $\pi_{\Delta}$. Given any $x_{t}\in\mathcal{X}$, the performance degeneration $I^{\text{direct}}$ of the directly forecasting belief $b_{\theta}$ can be bounded as follows respectively.
@@
It is obvious that the belief errors of recursively forecasting grow much faster than the directly forecasting, which is not strictly related to the delay length. We also empirically show that Proposition 5.8 is always held in Section 6.2.1. In the future, we will theoretically investigate the sample complexity of recursive and direct forecasting beliefs.
@@
Table 1: Performance on MuJoCo with Deterministic Delays. The best performance is underlined, and the best belief-based method is in red.
@@ 6.1 Experimental Setting
We adopt D4RL (Fu et al., 2020) and MuJoCo (Todorov et al., 2012) as the offline dataset and the benchmark, respectively, to evaluate our DFBT-SAC. For the baselines, we choose the SOTA augmentation-based methods (A-SAC (Haarnoja et al., 2018), BPQL (Kim et al., 2023), and ADRL (Wu et al., 2024b)) and belief-based methods (DATS (Chen et al., 2021a), D-Dreamer (Karamzade et al., 2024), and D-SAC (Liotet et al., 2021)). All of the belief-based baselines and our DFBT-SAC are trained on the same D4RL dataset. We first investigate the prediction accuracy of beliefs (Section 6.2.1), followed by the performance comparison with deterministic and stochastic delays (Section 6.2.2). Additionally, we conduct ablation studies on the multi-step bootstrapping of our DFBT-SAC (Section 6.2.3).
@@ 6.2.1 Belief Prediction Accuracy
We first evaluate the state prediction accuracy of our DFBT in D4RL offline datasets. Using the L1 norm as the belief error, we report the error curves increasing with delays of the belief of DATS, D-Dreamer, D-SAC and our DFBT-SAC in Figure 2. All methods are trained on the D4RL mixed dataset, including random, medium and expert policy demonstrations. The implementation details are provided in Appendix A. From the results, we can tell that our DFBT can address the compounding errors effectively via directly forecasting delayed observations, thus maintaining the best prediction accuracy with increased delays, which is consistent with our theoretical results. We further provide the belief qualitative comparison in Appendix E.
@@ 6.2.2 Performance Comparison
We report the normalized return $R_{\text{norm}}:=\frac{R_{\text{alg}}-R_{\text{random}}}{R_{\text{sac}}-R_{\text{random}}}$, where $R_{\text{alg}},R_{\text{sac}},R_{\text{random}}$ are the return of the algorithm, delay-free SAC, and random returns, respectively. Each method is evaluated across 5 random seeds, and the implementation details are provided in Appendix A. Additional experiment results and learning curves for different delays are provided in the Appendix C and Appendix D, respectively.
@@
The performance of DFBT-SAC and baselines is evaluated on MuJoCo with deterministic delays (8, 32, and 128) reported in Table 1, showing that our DFBT-SAC overall outperforms other belief-based methods significantly. Specifically, DFBT-SAC shows comparable performance with the SOTA augmentation-based methods (BPQL and ADRL) on tasks with relatively short delays (8). However, our DFBT-SAC yields a leading performance on challenging tasks when the delays increase to long delays (32 and 128). Additionally, we also compare DFBT-SAC(1), the single-step bootstrapping version of DFBT-SAC. The results imply that the multi-step bootstrapping technique can further improve the performance effectively, validating our statement in Section 4.2. We will investigate the bootstrapping steps in the later ablation study.
@@
We also evaluate DFBT-SAC on MuJoCo with stochastic delays which follow the uniform distribution $U$. As shown in Table 2, our DFBT-SAC remarkably outperforms all belief-based baselines on all tasks with all delay settings ($U(1,8)$, $U(1,32)$, and $U(1,128)$). Especially for $U(1,128)$ delays, DFBT-SAC also performs approximately $144\%$, $180\%$, and $100\%$ better than the second best baselines on HalfCheetah-v2, Hopper-v2, and Walker2d-v2, respectively.
@@
As summarised in Figure 3, we report the learning curves on MuJoCo with $128$ and $U(1,128)$ delays. Our DFBT-SAC exhibits a leading learning efficiency across all these challenging delays. Especially in the MuJoCo with 128 delays, DFBT-SAC can learn remarkably faster than all baselines, resulting in $173.3\%$ (HalfCheetah-v2), $122.2\%$ (Hopper-v2), and $263.6\%$ (Walker2d-v2) final performance improvement better than the second-best baselines.
@@
Based on the above experimental results, we can observe a general trend in both deterministic and stochastic delays that DFBT-SAC shows comparable performance with SOTA augmentation-based methods under short delays scenarios, in which augmentation-based methods generally have better performance than other belief-based approaches. Since delays are short, the dimensionality of the augmented state space is proper for learning, and without the approximation errors from belief representation. However, as delays increase in both scenarios, DFBT-SAC unleashes its advantages. This empirical finding validates our theoretical analysis in Section 5. For augmentation-based methods, as the delays increase, the augmented state in augmentation-based methods grows too rapidly, making efficient policy optimisation infeasible. For recursively forecasting belief methods, the compounding errors in belief function approximation grow exponentially with increasing delays, as demonstrated in Figure 2 and Section 6.2.1. This results in inaccurate predictions of delayed observations and undermines subsequent policy training. In contrast, DFBT’s prediction error remains largely unaffected by the length of delays, enabling DFBT-SAC to maintain strong performance even in scenarios with long delays.
@@ 6.3 Limitations and Challenges
Online Belief Learning
In this paper, we mainly consider learning DFBT from the offline dataset, which means the performance of the belief is subject to the quality and quantity of the dataset. However, directly learning belief in the online environment will introduce an auxiliary representation task, destabilising the learning process.
@@ Appendix B Theoretical Analysis
We use $b^{(i)}(\cdot|{x_{t}})(i=1,\cdots,\Delta)$ to note that the belief function with the specific delays $i$. For instance, $b^{(\Delta)}(\cdot|{x_{t}})=b(\cdot|{x_{t}})$ and $b^{(1)}(\cdot|{x_{t}})=\mathcal{P}(\cdot|{s_{t},a_{t}})$.
@@
For the delay-free policy $\pi$ and the delayed policy $\pi_{\Delta}$. Given any $x_{t}\in\mathcal{X}$, the performance degeneration $I^{\text{direct}}$ of the directly forecasting belief $b_{\theta}$ can be bounded as follows respectively.
@@
We report additional experimental results on MuJoCo including more different patterns of deterministic and stochastic delays in Table 6 and Table 7, respectively.
Table 6: Performance on MuJoCo with Deterministic Delays. The best performance is underlined, and the best belief-based method is in red.

### v2

@@ 4.2 Multi-step Bootstrapping with DFBT
Then, we present our practical delayed RL method, named DFBT-SAC by incorporating the trained DFBT $b_{\theta}$ with Soft Actor-Critic (Haarnoja et al., 2018). To improve the learning efficiency with the DFBT, the critic of DFBT-SAC is bootstrapped on the states predicted by the DFBT with the multi-step learning (Sutton & Barto, 2018; Hessel et al., 2018) and delay-free training techniques (Wu et al., 2024b; Kim et al., 2023). Specifically, given multi-step data $(x^{\text{tokens}}_{t},s_{t-\Delta+1:t})$, the critic $Q_{\psi}$ parameterized by $\psi$ is updated via:
@@ 5.2 Directly Forecasting Belief: Errors Analysis
Next, we theoretically show that our directly forecasting belief can effectively address the compounding errors. The sequence-modeling method estimates the states $s_{t-\Delta+1:t}$ based on the augmented state $x_{t}$ in parallel, instead of invoking $\Delta$ times iteratively, which effectively alleviates the source of compounding errors. Especially, the attention mechanism in the transformer can selectively capture the long-range relationships within the augmented state with long delays. We assume that the belief error between the directly forecasting belief and the ground-truth belief is bounded (5.6).
@@
The distance between the directly forecasting belief belief $b_{\theta}$ parameterized by $\theta$ and the ground-truth belief $b$ is bounded, it satisfies that $\forall x_{t}\in\mathcal{X}$, we have $\mathcal{W}(b(\cdot|{x_{t}})||b_{\theta}(\cdot|{x_{t}}))\leq\epsilon_{direct}$, where $\epsilon_{direct}:=\max_{i=1,\ldots,\Delta}\mathcal{W}(b^{(i)}(\cdot|{x_{t}})||b^{(i)}_{\theta}(\cdot|{x_{t}}))$.
@@
For the delay-free policy $\pi$ and the delayed policy $\pi_{\Delta}$. Given any $x_{t}\in\mathcal{X}$, the performance degeneration $I^{\text{direct}}$ of the directly forecasting belief $b_{\theta}$ can bounded as follows respectively.
@@
It is obvious that the belief errors of recursively forecasting grows much faster than the directly forecasting, which is not strictly related with the delay length. We also empirically show that Proposition 5.8 is always held in Section 6.2.1. In the future, we will theoretically investigate the sample complexity of recursive and direct forecasting beliefs.
@@
Table 1: Performance on MuJoCo with Deterministic Delays. The best performance is underlined, the best belief-based method is in red.
@@ 6.1 Experimental Setting
We adopt D4RL (Fu et al., 2020) and MuJoCo (Todorov et al., 2012) as the offline dataset and the benchmark respectively to evaluate our DFBT-SAC. For the baselines, we choose the SOTA augmentation-based methods (A-SAC (Haarnoja et al., 2018), BPQL (Kim et al., 2023), and ADRL (Wu et al., 2024b)) and belief-based methods (DATS (Chen et al., 2021a), D-Dreamer (Karamzade et al., 2024), and D-SAC (Liotet et al., 2021)). All of the belief-based baselines and our DFBT-SAC are trained on the same D4RL dataset. We first investigate the prediction accuracy of beliefs (Section 6.2.1), followed by the performance comparison with deterministic and stochastic delays (Section 6.2.2). Additionally, we conduct ablation studies on the multi-step bootstrapping of our DFBT-SAC (Section 6.2.3).
@@ 6.2.1 Belief Prediction Accuracy
We first evaluate the state prediction accuracy of our DFBT in D4RL offline datasets. Using the L1 norm as the belief error, We report the error curves increasing with delays of the belief of DATS, D-Dreamer, D-SAC and our DFBT-SAC in Figure 2. All methods are trained on the D4RL mixed dataset including random, medium and expert policy demonstrations. The implementation details are provided in Appendix A. From the results, we can tell that our DFBT can address the compounding errors effectively via directly forecasting delayed observations, thus maintaining the best prediction accuracy with increased delays, which is consistent with our theoretical results. We further provide the belief qualitative comparison in Appendix E.
@@ 6.2.2 Performance Comparison
We report the normalized return $R_{\text{norm}}:=\frac{R_{\text{alg}}-R_{\text{random}}}{R_{\text{sac}}-R_{\text{random}}}$, where $R_{\text{alg}},R_{\text{sac}},R_{\text{random}}$ are the return of the algorithm, delay-free SAC, and random returns, respectively. Each method is evaluated across 5 random seeds, and the implementation details are provided in the Appendix A. Additional experiment results and learning curves for different delays are provided in the Appendix C and Appendix D, respectively.
@@
The performance of DFBT-SAC and baselines are evaluated on MuJoCo with deterministic delays (8, 32, and 128) reported in Table 1, showing that our DFBT-SAC overall outperforms other belief-based methods significantly. Specifically, DFBT-SAC shows comparable performance with the SOTA augmentation-based methods (BPQL and ADRL) on tasks with relatively short delays (8). However, our DFBT-SAC yields a leading performance on challenging tasks when the delays increase to long delays (32 and 128). Additionally, we also compare DFBT-SAC(1), the single-step bootstrapping version of DFBT-SAC. The results imply that the multi-step bootstrapping technique can further improve the performance effectively, validating our statement in Section 4.2. We will investigate the bootstrapping steps in the later ablation study.
@@
We also evaluate DFBT-SAC on MuJoCo with stochastic delays which follow the uniform distribution $U$. As shown in Table 2, our DFBT-SAC remarkably outperforms all belief-based baselines on all tasks with all delays settings ($U(1,8)$, $U(1,32)$, and $U(1,128)$). Especially for $U(1,128)$ delays, DFBT-SAC also performs approximately $144\%$, $180\%$, and $100\%$ better than the second best baselines on HalfCheetah-v2, Hopper-v2, and Walker2d-v2, respectively.
@@
As summarized in Figure 3, we report the learning curves on MuJoCo with $128$ and $U(1,128)$ delays. Our DFBT-SAC exhibits a leading learning efficiency across all these challenging delays. Especially in the MuJoCo with 128 delays, DFBT-SAC can learn remarkably faster than all baselines, resulting in $173.3\%$ (HalfCheetah-v2), $122.2\%$ (Hopper-v2), and $263.6\%$ (Walker2d-v2) final performance improvement better than the second-best baselines.
@@
Based on the above experimental results, we can observe a general trend in both deterministic and stochastic delays that DFBT-SAC shows comparable performance with SOTA augmentation-based methods under short delays scenarios, in which augmentation-based methods generally have better performance than other belief-based approaches. Since when delays are short, the dimensionality of the augmented state space is proper for learning, and without the approximation errors from belief representation. However, as delays increase in both scenarios, DFBT-SAC unleashes its advantages. This empirical finding validates our theoretical analysis in Section 5. For augmentation-based methods, as the delays increase, the augmented state in augmentation-based methods grows too rapidly, making efficient policy optimization infeasible. For recursively forecasting belief methods, the compounding errors in belief function approximation grow exponentially with increasing delays, as demonstrated in Figure 2 and Section 6.2.1. This results in inaccurate predictions of delayed observations and undermines subsequent policy training. In contrast, DFBT’s prediction error remains largely unaffected by the length of delays, enabling DFBT-SAC to maintain strong performance even in scenarios with long delays.
@@ 6.2.3 Ablation Study on Bootstrapping Steps
Table 4: Performance comparison of DFBT-SAC with different belief training methods on MuJoCo tasks with 32 delays. The best performance is underlined.
Task Online Offline Offline + Fine-tuning
HalfCheetah-v2 $0.11\pm 0.39$ $0.42\pm 0.12$ $0.39\pm 0.04$
Hopper-v2 $0.10\pm 0.58$ $0.68\pm 0.20$ $0.84\pm 0.04$
Walker2d-v2 $0.09\pm 0.27$ $0.64\pm 0.10$ $0.96\pm 0.32$
6.2.4 Evaluation on Additional MuJoCo tasks.
we conducted experiments on the Pusher-v2, Reacher-v2, and Swimmer-v2 with deterministic 32 delays. The offline datasets (500k samples) are collected from a SAC policy, with other settings unchanged. The results are shown in Table 5, showing that DFBT-SAC achieves superior performance on these tasks.
Table 5: Performance comparison on additional MuJoCo tasks with 32 deterministic delays. The best performance is underlined.
Task A-SAC BPQL ADRL DATS D-Dreamer D-SAC DFBT-SAC (ours)
Pusher-v2 $0.05\pm 0.00$ $0.93\pm 0.58$ $0.95\pm 0.24$ $0.92\pm 0.10$ $0.87\pm 0.18$ $0.94\pm 0.04$ $1.04\pm 0.24$
Reacher-v2 $0.89\pm 0.08$ $0.83\pm 0.06$ $0.85\pm 0.01$ $0.82\pm 0.13$ $0.84\pm 0.02$ $0.88\pm 0.07$ $0.93\pm 0.06$
Swimmer-v2 $0.27\pm 0.05$ $0.80\pm 0.14$ $0.60\pm 0.06$ $0.25\pm 0.05$ $0.21\pm 0.07$ $0.30\pm 0.07$ $1.01\pm 0.27$
Table 6: Performance on stochastic MuJoCo with deterministic 128 delays. The best performance is underlined.
Task A-SAC BPQL ADRL DATS D-Dreamer D-SAC DFBT-SAC (ours)
HalfCheetah-v2 $0.00\pm 0.03$ $0.01\pm 0.05$ $0.13\pm 0.04$ $0.13\pm 0.03$ $0.07\pm 0.01$ $0.00\pm 0.04$ $0.35\pm 0.04$
Hopper-v2 $0.03\pm 0.04$ $0.08\pm 0.05$ $0.06\pm 0.04$ $0.05\pm 0.06$ $0.04\pm 0.05$ $0.04\pm 0.05$ $0.13\pm 0.22$
Walker2d-v2 $0.06\pm 0.02$ $0.04\pm 0.01$ $0.08\pm 0.01$ $0.06\pm 0.03$ $0.10\pm 0.03$ $0.05\pm 0.02$ $0.30\pm 0.07$
6.2.5 Ablation Results of Belief Training.
Some task-specific information may be missing if the belief is frozen in the RL process, leading to limited performance improvement. This issue can be mitigated by fine-tuning the belief within the RL process. The results, shown in Table 4, demonstrate that fine-tuning helps the DFBT capture the task-specific information with better performance. Note that there are many potential methods for capturing task-specific information, not limited to fine-tuning DFBT. Belief learning from scratch in the online RL process always suffers from instability issues. Therefore, in this paper, we separate belief learning from the online RL process, which allows us to investigate the belief component solely, eliminate potential influences from the RL side.
6.2.6 Stochastic MuJoCo.
We conducted additional experiments on the stochastic MuJoCo tasks with a probability of 0.001 for the unaware noise and deterministic 128 delays. As shown in Table 6, the results demonstrate that our DFBT-SAC achieves superior performance in these stochastic MuJoCo tasks.
6.2.7 Inference Speed Comparison.
We conducted additional experiments on computational efficiency. As shown in Table 7, the results demonstrate that directly forecasting belief maintains a consistent and stable inference speed (around 4 ms) across different delays. In contrast, the recursively forecasting belief experiences inference speed issues as delays increase. In HalfCheetah-v2 with 128 delays, the training times of DATS and D-Dreamer are around 10 hours and 15 hours, respectively, while those of D-SAC and DFBT-SAC are both around 6 hours.
Table 7: Inference speed (ms) comparison in HalfCheetah-v2.
Delays DATS D-Dreamer D-SAC DFBT-SAC
8 $1.10\pm 0.02$ $1.85\pm 0.01$ $4.03\pm 0.03$ $4.18\pm 0.04$
32 $3.85\pm 0.06$ $6.80\pm 0.04$ $4.03\pm 0.04$ $4.18\pm 0.04$
128 $14.97\pm 0.22$ $26.51\pm 0.19$ $4.03\pm 0.03$ $4.15\pm 0.05$
@@ 6.3 Limitations and Challenges
Online Belief Learning.
In this paper, we mainly consider learning DFBT from the offline dataset, which means the performance of the belief is subject to the quality and quantity of the dataset. However, direct learning belief in the online environment will introduce an auxiliary representation task, destabilizing the learning process.
@@ 7 Conclusion
Acknowledgement
We sincerely appreciate the insightful suggestions provided by ICML 2025 reviewers, which help us improve the quality of this paper. We sincerely acknowledge the support by the grant EP/Y002644/1 under the EPSRC ECR International Collaboration Grants program, funded by the International Science Partnerships Fund (ISPF) and the UK Research and Innovation. We also sincerely acknowledge the support by Taiwan NSTC under Grant Number NSTC-112-2221-E-002-168-MY3, and from King Abdullah University of Science and Technology (KAUST) - Center of Excellence for Generative AI, under award number 5940. Simon Sinong Zhan and Qi Zhu are partially supported by the US National Science Foundation grants 2324936 and 2328973.
Impact Statement
This paper aims to advance the field of Machine Learning. While acknowledging there are many potential societal consequences of our work, we believe that none need to be specially highlighted here.
@@ Appendix B Theoretical Analysis
We use $b^{(i)}(\cdot|{x_{t}})(i=1,\ldots,\Delta)$ to note that the belief function with the specific delays $i$. For instance, $b^{(\Delta)}(\cdot|{x_{t}})=b(\cdot|{x_{t}})$ and $b^{(1)}(\cdot|{x_{t}})=\mathcal{P}(\cdot|{s_{t},a_{t}})$.
@@
For the delay-free policy $\pi$ and the delayed policy $\pi_{\Delta}$. Given any $x_{t}\in\mathcal{X}$, the performance degeneration $I^{\text{direct}}$ of the directly forecasting belief $b_{\theta}$ can bounded as follows respectively.
@@
We report additional experimental results on MuJoCo, including more different patterns of deterministic and stochastic delays in Table 10 and Table 11, respectively.
Table 10: Performance on MuJoCo with Deterministic Delays. The best performance is underlined, the best belief-based method is in red.